Shares in Lex plc are currently trading at £5.80 per share with an equity beta of 1.32.

The current equity market premium and the risk-free rate of return are 6% and 1.35% respectively.

In the last financial year Lex reported earnings before interest and tax (EBIT) of £240m and an interest cover ratio of 3.2. Lex’s outstanding debt finance totals £1200m

The current capital structure of Lex comprises 45% debt and 55% equity.

The corporate tax rate is 20.5%

Required:

Calculate the cost of equity, the cost of debt and the weighted average cost of capital (WACC) for Lex plc and explain your calculations
Answer #1

Final answers .....

Cost of equity = 9.27 %

Cost of debt = 4.97%

WACC = 7.33 %

Explanation

Cost of equity = Rf + Beta * ( Rm - Rf )

Rf = Risk free rate = 1.35

beta = 1.32

Market premium = Rm - Rf = 6

= 1.35 + 1.32 * ( 6 )

= 9.27 %

Cost of Debt.

Interest = EBIT / Interest coverage ratio = 240 / 3.20 = $ 75 millions

Cost of debt ( before tax ) = 75 / 1200 * 100 = 6.25 %

Cost of debt ( after tax ) = 6.25 * ( 1 - tax rate ) = 6.25 * (1 - 0.205) = 4.96875 % ............. or 4.97 %

WACC

= Cost of equity * ( weight of equity ) + Cost of debt * ( weight of Debt )

= 9.27 * (0.55) + 4.96875 * (0.45 ) = 7.33 %

Kindly rate my answer.